# Continuous Ensemble Weather Forecasting with Diffusion models

### v1

@@ 4.3 Autoregressive roll-outs with continuous interpolation
Continuous forecasting is effective for forecasting hours to days but can struggle to forecast longer lead times where the correlation is weaker. Iterative forecasting excels at long lead times when used with longer (24h) timesteps (Bi et al., 2023), but comes at the loss of temporal resolution. In our framework, it becomes possible to sample both iterative and continuous forecasts with the same model. We propose to leverage this by iterating on a longer timestep and forecasting the intermediate timesteps using Continuous Ensemble Forecasting, as outlined in Alg. 3. We refer to this combined method as Autoregressive Rollouts with Continuous Interpolation (ARCI). Our method limits the error accumulation without sacrificing temporal resolution. This allows for producing forecasts at an arbitrary fine temporal resolution, while retaining the accuracy of the best iterative methods throughout the whole forecast. By limiting the number of iterative steps, more of the forecast also becomes parallelizable, allowing for rapidly generating forecasts on large compute clusters. Furthermore, we can straightforwardly use different time resolutions during different parts of the forecast trajectories, by for instance forecasting with 1h steps for the first few days and then switching to longer time steps for long lead times. Note that ARCI can be used with either fixed (with Alg. 1) or stochastic (with Alg. 2) driving noise. However, we emphasize that these two algorithms are probabilistically equivalent for all time marginals, and only differ in the autocorrelation of forecast trajectories.
@@
1:input: Initial conditions $x_{-M:0}$, interpolation times $\{t_{i}\}_{i=1}^{N}$, ensemble size $n_{\text{ens}}$, autoregressive steps $J$, network $S_{\theta}$
2:for $j=0$ to $J-1$ do
3:     $\{x^{k}_{jN+i}\}^{k=1:n_{\text{ens}}}_{i=1:N}\leftarrow$ Alg. 1$(x_{jN{-}M{:}jN},\{t_{i}\}_{i=1}^{N},n_{\text{ens}},S_{\theta})$triangleright Also possible to use Alg. 2 return $\{x^{k}_{i}\}^{k=1:n_{\text{ens}}}_{i=1:JN}$
@@ 5 Experiments
We propose to use the ARCI model described in algorithm 3 referred to as ARCI-24/6h. We train it to forecast $t\in\{6,12,18,24\}$ (hours) and roll it out autoregressively with 24 h steps, hence the name. Training is done on a 40GB NVIDIA A100 GPU and takes roughly 2 days. We emphasize again that the choice of algorithm inside ARCI results in probabilistically equivalent forecasts for all time marginals, and only differ in the autocorrelation of forecast trajectories. Hence, for all metrics below that are computed for specific lead times we only report results for one version of the algorithm. We return to the difference between Alg. 1 and Alg. 2 when studying the temporal difference below.
To evaluate the effectiveness of our approach, we compare it to other MLWP baselines. Deterministic is a deterministic model trained using MSE-loss on a single 6 hour time step, and unrolled up to 10 days. AR-6/24h is a diffusion model trained only on forecasting a single fixed $\delta$ ahead, and then autoregressively unrolled up to 10 days. This is the forecasting setup of (Price et al., 2024). CI-6h is a diffusion model performing continuous forecasting conditioned on a specific lead time. It is trained on uniformly sampled lead times from $\{k\delta\}_{k=1}^{40}$, with $\delta=\text{6 h}$. This is the method proposed in alg. 1. Sampling a 10-day forecast with 6h resolution for a single member from AR-6h takes 32 seconds, but by parallelizing the 6h timesteps in ARCI-24/6h this reduces to 8 seconds.
To compare against another family of ensemble forecasting models from the literature we retrain the Graph-EFM model (Oskarsson et al., 2024) on our exact data setup. Graph-EFM is a graph-based latent-variable model that produces forecasts by 6 h iterative rollout steps. For all models, unless otherwise specified, we condition on the two previous timesteps $\Omega=\{0,-\delta\}$ and sample 50 ensemble members at each initialization time. All models except Graph-EFM use the same architecture based on the U-net in Karras et al. (2022) as presented in appendix B.
@@ 5.1 Results
Table 1 and figure 3 show metrics for a selection of lead times and variables. Scores for the remaining variables are listed in appendix C. All probabilistic models show a clear improvement over the deterministic model. CI-6h performs well on short-term forecasting but struggles at longer horizons. This is likely due to the challenge of learning any useful relationships between initial states and later lead times, which are weakly correlated. ARCI-24+6h outperforms all models at 6h resolution, including the external baseline Graph-EFM, and matches the best overall model AR-24h in almost all scores. All diffusion-based models have SSR $<$ 1, indicating some systematic underdispersion.
Table 1: Selection of results for 5 and 10 day forecasting of geopotential at 500 hPa (z500) and temperature at 850 hPa (t850). For RMSE and CRPS, lower values are better, and SSR should be close to 1. The best values are marked with bold and second best underlined.
@@
Figure 3: RMSE, CRPS and SSR for temperature at 850 hPa (t850) with a selection of models at 6h resolution.
Qualitative results.
@@
Temporal Difference.
Autoregressively sampled forecast trajectories are necessarily continuous. Since there is no standard way of measuring the continuity of a forecast, we propose using the mean temporal difference ${\Delta X=|X(t)-X(t-1)|}$ as a measure of forecast continuity. Figure 5 shows $\Delta X$ for a CI-1h continuous model trained up to 24 hours with 1-hour timesteps. Compared to using different noise at each step ($\rho\to\infty)$, the temporal difference of our model $(\rho=0)$ stays close to the temporal difference of the data. This supports our claim that continuous ensemble forecasting produces continuous trajectories.
Figure 5 also shows the temporal difference of the continuous model above. When the noise is fixed ($\rho=0$) the temporal difference decreases with lead time, corresponding to predictions with smaller temporal variations. Letting the noise vary with noise factor $(\rho=\text{ln}10)$ as in alg. 2 stops this from happening. The bias between $\Delta X$ of the data and our model is likely due to the model producing slightly blurrier forecasts, making the differences smaller.
@@
Our ARCI method allows for producing forecasts at arbitrary fine temporal resolution while retaining the accuracy of methods taking longer autoregressive steps. We here demonstrate this by producing hourly forecasts. Figure 6 shows the scores of a selection of models for z500 for 10-day forecasts at 1h resolution. The AR-1h model has the same setup as AR-6h but on 1h resolution. ARCI-24/1h and ARCI-24/2h* are both trained with $\Omega=\{0,-24\}$ to have access to the same information at each timestep. The autoregressive AR-1h model performs much worse than on 6 or 24 hours. The continuous model, however, does not lose performance by increasing the temporal resolution, making the 1h timestep forecasts as skillful as the 24 hour ones. An alternative to directly producing forecasts at a fine temporal resolution would be to linearly interpolate the forecasts sampled using an autoregressive model. In fig. 9 in appendix C we show that linearly interpolated forecasts behave much worse than the ARCI model on both 1 and 6-hour resolution.
@@
Figure 6: RMSE, CRPS, and SSR for geopotential at 500 hPa (z500) with a selection of models at 1h resolution.
@@ 6 Conclusion
We present Continuous Ensemble Forecasting, a novel framework for probabilistic MLWP that increases the efficiency, accuracy, and flexibility of weather forecasts at high temporal resolution. When combined with autoregressive prediction, our ARCI method produces 10-day forecasts with a 1-hour resolution that matches the accuracy of a purely autoregressive model with 24-hour steps. With this work, we hope to show that the possibilities with generative modeling for spatio-temporal predictions are still largely unexplored and a fruitful area of research.
@@
While our proposed framework achieves good results on 5.625°Weatherbench (Rasp et al., 2020) data, we have yet to show that the method scales to problems with higher spatial resolution. Additionally, as the lead time increases, the correlation between initial and future states becomes weaker, limiting the application of continuous forecasting. While our method parallelizes more of the sampling than previous autoregressive models, solving the probability flow ODE in eq. 2 still requires many sequential forward passes through the network. Sampling is thus still slower than for latent variable models, but the predicted distribution more accurate. Finally, due to computational limitations, we were not able to retrain our models several times and as such do not report error bars in our results.
@@ Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. Our computations were enabled by the Berzelius resource at the National Supercomputer Centre, provided by the Knut and Alice Wallenberg Foundation.
@@ Appendix B Model
$D_{\theta}(X_{t}^{\sigma};\sigma,X(\Omega),t)=c_{\text{skip}}(\sigma)\cdot X_{t}^{\sigma}+c_{\text{out}}(\sigma)\cdot F_{\theta}\left(c_{\text{in}}(\sigma)\cdot X_{t}^{\sigma};c_{\text{noise}}(\sigma),X(\Omega),t\right),$
where $X_{t}^{\sigma}$ denotes a noisy version of the target time $X({t})$ at noise level $\sigma$, and $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$ and $c_{\text{noise}}$ are scaling functions taken from (Karras et al., 2022) defined in Tab. 2. These scaling functions $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$ and $c_{\text{noise}}$, are specifically chosen to handle the influence of the noise level within the network, allowing $D_{\theta}$ to adapt dynamically to different noise intensities without the need for adjusting the scale of $\sigma$ externally. Consequently, for consistency with the normalization of the data where $\sigma_{\text{data}}$ is set to 1, the lead time $t$ is also scaled to fit within the range $[0,1]$. This normalization ensures that the network inputs are uniformly scaled, enhancing the efficiency and effectiveness of the denoising process.
@@
The backbone of the diffusion model is a U-Net architecture. Our model is based on the one used in (Karras et al., 2022), reconfigured for our purposes with 32 filters as the base multiplier. It is built up by blocks configured as in fig. 7. The blocks consist of two convolutional layers and are constructed as in fig. 8. If the block is a down-/up-sample or if the number of input filters is different from the number of output filters, there is an additional skip layer from the input to the output. The time/noise embedding is fed directly into each block and not passed through the network. Unlike the network in Karras et al. (2022), our convolutions uses zero padding at the poles and periodic padding at the left/right edges. This periodic padding ensures periodicity over longitudes. The model has 3.5M parameters.
@@
$\text{d}\bm{z}=-\dot{\sigma}(s)\sigma(s)\nabla_{\bm{z}}\log p_{s}\left(\bm{z}\right)\text{d}s.$ (7)
We employ the second-order Heun’s method, a deterministic ODE solver, as outlined in Algorithm 4. For the noise parameters, we define the noise level function as $\sigma(s)=s$. Additionally, we set a noise level schedule to lower the noise during sampling from $\sigma_{\max}$ to $\sigma_{\min}$ over $N$ steps:
@@
1:procedure HeunSampler($D_{\theta}(\bm{z};\sigma,X(\Omega),t)$, $s_{i\in\{0,...,N\}}$, $Z$)
2:$\bm{z}_{0}\leftarrow\sigma^{2}(s_{0})\cdot Z$triangleright Generate initial sample at $s_{0}$
@@
4:     $\bm{d}_{i}\leftarrow\frac{\dot{\sigma}(s_{i})}{\sigma(s_{i})}(\bm{z}_{i}-D_{\theta}(\bm{z}_{i};\sigma(s_{i}),X(\Omega),t))$triangleright Evaluate $d\bm{z}/ds$ at $s_{i}$
5:     $\bm{z}_{i+1}\leftarrow\bm{z}_{i}+(s_{i+1}-s_{i})\bm{d}_{i}$triangleright Take Euler step from $s_{i}$ to $s_{i+1}$
@@
7:         $\bm{d}_{i}^{\prime}\leftarrow{\frac{\dot{\sigma}(s_{i+1})}{\sigma(s_{i+1})}}(\bm{z}_{i+1}-D_{\theta}(\bm{z}_{i+1};\sigma(s_{i+1}),X(\Omega),t))$triangleright Evaluate $d\bm{z}/ds$ at $s_{i+1}$
8:         $\bm{z}_{i+1}\leftarrow\bm{z}_{i}+\frac{1}{2}(s_{i+1}-s_{i})\left(\bm{d}_{i}+\bm{d}_{i}^{\prime}\right)$triangleright Explicit trapezoidal rule at $s_{i+1}$      return $\bm{z}_{N}$triangleright Return noise-free sample at $s_{N}$

### v2

@@ 4.3 Autoregressive roll-outs with continuous interpolation
Continuous forecasting is effective for forecasting hours to days but can struggle to forecast longer lead times where the correlation is weaker. Autoregressive forecasting excels at long lead times when used with longer (24h) timesteps (Bi et al., 2023), but comes at the loss of temporal resolution. In our framework, it becomes possible to sample both autoregressive and continuous forecasts with the same model. We propose to leverage this by iterating on a longer timestep and forecasting the intermediate timesteps using Continuous Ensemble Forecasting, as outlined in Alg. 3. We refer to this combined method as Autoregressive Rollouts with Continuous Interpolation (ARCI). Our method limits the error accumulation without sacrificing temporal resolution. This allows for producing forecasts at an arbitrary fine temporal resolution, while retaining the accuracy of the best autoregressive methods throughout the whole forecast. By limiting the number of autoregressive steps, more of the forecast also becomes parallelizable, allowing for rapidly generating forecasts on large compute clusters. Furthermore, we can straightforwardly use different time resolutions during different parts of the forecast trajectories, by for instance forecasting with 1h steps for the first few days and then switching to longer timesteps for long lead times. Note that ARCI can be used with either fixed (with Alg. 1) or stochastic (with Alg. 2) driving noise. However, we emphasize that these two algorithms are probabilistically equivalent for all time marginals, and only differ in the autocorrelation of forecast trajectories.
@@
1:input: Initial conditions $x_{-L:0}$, interpolation times $\{t_{i}\}_{i=1}^{N}$, ensemble size $n_{\text{ens}}$, autoregressive steps $M$, network $S_{\theta}$
2:for $m=0$ to $M-1$ do
3:     $\{x^{j}_{mN+i}\}^{j=1:n_{\text{ens}}}_{i=1:N}\leftarrow$ Alg. 1$(x_{mN{-}L{:}mN},\{t_{i}\}_{i=1}^{N},n_{\text{ens}},S_{\theta})$triangleright Also possible to use Alg. 2 return $\{x^{j}_{i}\}^{j=1:n_{\text{ens}}}_{i=1:MN}$
@@ 5 Experiments
We propose to use the ARCI model described in algorithm 3 referred to as ARCI-24/6h. We train it to forecast $t\in\{6,12,18,24\}$ (hours) and roll it out autoregressively with 24h steps, hence the name. Training is done on a 40GB NVIDIA A100 GPU and takes roughly 2 days. We emphasize again that using fixed, correlated or uncorrelated noise results in probabilistically equivalent forecasts for all time marginals, and only differ in the autocorrelation of forecast trajectories. Hence the choice of algorithm inside ARCI does not matter, and for all metrics below that are computed for specific lead times we only report results for one version of the algorithm. We return to the difference between Alg. 1 and Alg. 2 when studying the temporal difference below.
[…]
To compare against another family of ensemble forecasting models from the literature we retrain the Graph-EFM model (Oskarsson et al., 2024) on our exact data setup. Graph-EFM is a graph-based latent-variable model that produces forecasts by 6h iterative rollout steps. We also present new results from the DYffusion model Rühling Cachay et al. (2023), trained on ERA5 data at 1h resolution. All models except Graph-EFM use the same U-net with 3.5M parameters. The architecture is based on the U-net in Karras et al. (2022) and presented in appendix B. Due to its specific forward-backward process, DYffusion only supports conditioning on the current timestep, $\Omega=\{0\}$. For all other models we condition on the two previous timesteps $\Omega=\{0,-\delta\}$.
@@ 5.1 Results
Table 1 and figure 4(a) show metrics for a selection of lead times and variables. Scores for the remaining variables are listed in appendix C. For probabilistic models on 6h resolution, we sample 50 ensemble members at each initialization time. All probabilistic models show a clear improvement over the deterministic model. CI-6h performs well on short-term forecasting but struggles at longer horizons. This is likely due to the challenge of learning any useful relationships between initial states and later lead times, which are weakly correlated. ARCI-24+6h outperforms all models at 6h resolution, including the external baseline Graph-EFM and the GenCast setup AR-6h, and matches AR-24h in almost all scores. All diffusion-based models have SSR $<$ 1, indicating some systematic underdispersion. In tables 6, 7 in appendix C we present error bars calculated for the ARCI-24/6h model, which shows that the model is robust to network initialization.
Table 1: Selection of results for 5 and 10 day forecasting using models with 6h resolution for geopotential at 500 hPa (z500) and temperature at 850 hPa (t850). For RMSE and CRPS, lower values are better, and SSR should be close to 1. The best values are marked with bold and second best underlined. The AR-24h model is included for reference, but is not considered for best model since it operates at a coarser 24h resolution.
@@
AR-24h 2.55 1.24 0.89 3.25 1.6 0.96
(a) Selection of models at 6h resolution.
(b) Selection of models at 1h resolution.
Figure 4: RMSE, CRPS and SSR for temperature at 850 hPa (t850).
High Temporal Resolution.
Our ARCI method allows for producing forecasts at high temporal resolution while retaining the accuracy of methods taking longer autoregressive steps. We here demonstrate this by producing hourly forecasts. Figure 4(b) shows the scores of a selection of models for t850 for 10-day forecasts at 1h resolution. Since the 1h forecasts requires more computational effort, we use only 10 ensemble members instead of 50. The AR-1h model has the same GenCast-setup as AR-6h but on 1h resolution and ARCI-24/1h is trained in the same way as ARCI-24/6h but at 1h resolution and with $\Omega=\{0,-24\}$. The autoregressive AR-1h model performs much worse than on 6 or 24 hours. The continuous model, however, does not lose performance by increasing the temporal resolution, making the 1h timestep forecasts as skillful as the 24 hour ones. Although DYffusion achieves much better results than AR-1h, it is beaten by ARCI-24/1h in all variables and lead times.
Forecast continuity.
@@
By correlating the noise across timesteps, our continuous ensemble forecasting method can produce temporally consistent ensemble trajectories without resorting to autoregressive predictions. Here we give empirical justification for this claim. Since autoregressively sampled forecast trajectories are necessarily continuous, there is no standard way of measuring the continuity of a forecast. We propose using the mean temporal difference ${\Delta X=|X(t)-X(t-1)|}$ as a measure of forecast continuity. Figure 5 shows $\Delta X$ for a CI-1h continuous model trained up to 24 hours with 1-hour timesteps. Compared to using different noise at each step ($\rho\to\infty)$, the temporal difference of our model $(\rho=0)$ stays close to the temporal difference of the data. This supports our claim that continuous ensemble forecasting produces continuous trajectories. When the noise is fixed ($\rho=0$) the temporal difference decreases with lead time, corresponding to predictions with smaller temporal variations. Letting the noise vary with noise factor $(\rho=\text{ln}10)$ as in alg. 2 stops this from happening. The bias between $\Delta X$ of the data and our model is likely due to the model producing slightly blurrier forecasts, making the differences smaller.
@@
One of the key advantages of continuous ensemble forecasting is its ability to generate forecasts at arbitrarily fine temporal resolutions. To test this, we train a model at a coarser temporal resolution and evaluate it at a higher resolution. Specifically, we consider the ARCI-24/2h* model, which is trained on lead times spaced 2 hours apart but evaluated at every 1-hour timestep. To ensure a fair comparison with ARCI-24/1h, ARCI-24/2h* is trained with $\Omega=\{0,-24\}$, providing access to the same information at each timestep. As shown in fig. 4(b), ARCI-24/2h* performs similarly to ARCI-24/1h, demonstrating its ability to generalize beyond the temporal resolution seen during training.
@@
An alternative to directly producing forecasts at a fine temporal resolution would be to linearly interpolate the forecasts sampled using an autoregressive model. In fig. 8 in appendix C we show that linearly interpolated forecasts behave much worse than the ARCI model on both 1 and 6-hour resolution.
@@ 6 Conclusion
We present Continuous Ensemble Forecasting, a novel framework for probabilistic MLWP that increases the efficiency, accuracy, and flexibility of weather forecasts at high temporal resolution. When combined with autoregressive prediction, our ARCI method can produce accurate 10-day global ensemble weather forecasts with a 1-hour resolution. It achieves the same accuracy as a purely autoregressive model with 24-hour steps and surpasses models like GenCast and DYffusion when reimplemented using the same neural network architecture. With this work, we hope to show that the possibilities with generative modeling for spatio-temporal predictions are still largely unexplored and a fruitful area of research.
@@
While our proposed framework achieves good results on 5.625°Weatherbench (Rasp et al., 2020) data, we have yet to show that the method scales to problems with higher spatial resolution. Additionally, as the lead time increases, the correlation between initial and future states becomes weaker, limiting the application of continuous forecasting. While our method parallelizes more of the sampling than previous autoregressive models, sampling from the diffusion model still requires many sequential forward passes through the network. Sampling is thus still slower than for latent variable models, but the predicted distribution more accurate.
@@ Acknowledgments
This work was financially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, the Swedish Research Council (project no: 2020-04122, 2024-05011), and the Excellence Center at Linköping–Lund in Information Technology (ELLIIT). Our computations were enabled by the Berzelius resource at the National Supercomputer Centre, provided by the Knut and Alice Wallenberg Foundation.
@@ Appendix B Model
Score-based generative models uses a parameterized score $S_{\theta}$ to sample from the target distribution. In practice, it turns out to be easier to learn a denoising network $D_{\theta}$ using the denoising score matching objective Ho et al. (2020). By a result shown in Vincent (2011), the score can then be retrieved using $S_{\theta}={(D_{\theta}-{z})}/{\sigma^{2}}$.
@@
$D_{\theta}(z,\sigma;X(\Omega),t)=c_{\text{skip}}(\sigma)\cdot z+c_{\text{out}}(\sigma)\cdot F_{\theta}\left(c_{\text{in}}(\sigma)\cdot z,c_{\text{noise}}(\sigma);X(\Omega),t\right),$
where $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$ and $c_{\text{noise}}$ are scaling functions taken from (Karras et al., 2022) defined in Tab. 2. These scaling functions $c_{\text{skip}}$, $c_{\text{out}}$, $c_{\text{in}}$ and $c_{\text{noise}}$, are specifically chosen to handle the influence of the noise level within the network, allowing $D_{\theta}$ to adapt dynamically to different noise intensities without the need for adjusting the scale of $\sigma$ externally. Consequently, for consistency with the normalization of the data where $\sigma_{\text{data}}$ is set to 1, the lead time $t$ is also scaled to fit within the range $[0,1]$. This normalization ensures that the network inputs are uniformly scaled, enhancing the efficiency and effectiveness of the denoising process.
@@
The backbone of the diffusion model is a U-Net architecture. Our model is based on the one used in (Karras et al., 2022), reconfigured for our purposes with 32 filters as the base multiplier. It is built up by blocks configured as in fig. 6. The blocks consist of two convolutional layers and are constructed as in fig. 7. If the block is a down-/up-sample or if the number of input filters is different from the number of output filters, there is an additional skip layer from the input to the output. The blocks at 16$\times$32 resolution additionally has attention with a single head. The time/noise embedding is fed directly into each block and not passed through the network. Unlike the network in Karras et al. (2022), our convolutions uses zero padding at the poles and periodic padding at the left/right edges. This periodic padding ensures periodicity over longitudes. The model has 3.5M parameters.
@@
$\mathrm{d}{z}=-\dot{\sigma}_{s}\sigma_{s}\nabla_{{z}}\log p_{s}\left({z}\right)\mathrm{d}s.$ (7)
We employ the second-order Heun’s method, a deterministic ODE solver, as outlined in Algorithm 4. For the noise parameters, we define the noise level function as $\sigma_{s}=s$. Additionally, we set a noise level schedule to lower the noise during sampling from $\sigma_{\max}$ to $\sigma_{\min}$ over $N$ steps:
@@
1:procedure HeunSampler($D_{\theta}({z};\sigma,X(\Omega),t)$, $s_{i\in\{0,...,N\}}$, $Z$)
2:${z}_{0}\leftarrow\sigma^{2}(s_{0})\cdot Z$triangleright Generate initial sample at $s_{0}$
@@
4:     $\bm{d}_{i}\leftarrow\frac{\dot{\sigma}_{s_{i}}}{\sigma_{s_{i}}}({z}_{i}-D_{\theta}({z}_{i};\sigma_{s_{i}},X(\Omega),t))$triangleright Evaluate $d{z}/ds$ at $s_{i}$
5:     ${z}_{i+1}\leftarrow{z}_{i}+(s_{i+1}-s_{i})\bm{d}_{i}$triangleright Take Euler step from $s_{i}$ to $s_{i+1}$
@@
7:         $\bm{d}_{i}^{\prime}\leftarrow{\frac{\dot{\sigma}_{s_{i+1}}}{\sigma_{s_{i+1}}}}({z}_{i+1}-D_{\theta}({z}_{i+1};\sigma_{s_{i+1}},X(\Omega),t))$triangleright Evaluate $d{z}/ds$ at $s_{i+1}$
8:         ${z}_{i+1}\leftarrow{z}_{i}+\frac{1}{2}(s_{i+1}-s_{i})\left(\bm{d}_{i}+\bm{d}_{i}^{\prime}\right)$triangleright Explicit trapezoidal rule at $s_{i+1}$      return ${z}_{N}$triangleright Return noise-free sample at $s_{N}$
@@
DYffusion Baseline
We have adapted the DYffusion model to work with weather forecasting using ERA5 data. This application was not considered in the original DYffusion papers which focused on sea-surface forecasting Rühling Cachay et al. (2023) and climate modeling Rühling Cachay et al. (2025). To ensure a fair evaluation, we used the same backbone U-Net as our diffusion models, and a lat-lon weighted mean-squared-error loss. We have trained it with a horizon of 24h, with 1h timesteps, making it directly comparable to ARCI-24/1h. The dropout rate was set to 0.2 and we used no artificial diffusion steps. For the training hyperparameters we used values similar to those proposed in the original paper.